14. Alpha Division of the Grande Company produces and sells two products. Each product's operating income and average capital resources are shown below: Product A: Operating Income $500,000; Average capital $5,000,000. Product B: Operating Income $350,000; Average capital $4,100,000. Assuming Alpha's manager has an opportunity to undertake an investment that would require a $500,000 investment and yield $40,000 in net operating income for its product B, what would the ROI be for the entire division? (Round your percentage answer to the nearest whole number.)

13.Alpha Division of the Grande Company produces and sells two products. Each product's operating income and average capital resources are shown below:

Product A: Operating Income $500,000; Average capital $5,000,000.

Product B: Operating Income $350,000; Average capital $4,100,000.

What is Alpha's ROI for product B? (Round your percentage answer to the nearest whole number.)
Answer #1

14) Calculate following

Product A Product B
Operating income 500000 350000+40000 = 390000
Average capital 5000000 4100000+500000 = 4600000

ROI = 890000/9600000 = 9%

13) ROI = 350000/4100000 = 9%